# Cone Ranking for Multi-Criteria Decision Makingthanks: Paper accepted and presented by D. Kostner at MODeM Workshop at ECAI 2023, Krakow, Poland

### 04 Dec 2023 v1

@@ Abstract
Recently introduced cone distribution functions from statistics are turned into multi-criteria decision making (MCDM) tools. It is demonstrated that this procedure can be considered as an upgrade of the weighted sum scalarization insofar as it absorbs a whole collection of weighted sum scalarizations at once instead of fixing a particular one in advance. Moreover, situations are characterized in which different types of rank reversal occur, and it is explained why this might even be useful for analyzing the ranking procedure. A few examples will be discussed and a potential application in machine learning is outlined.
@@ 1 The ranking problem
The main MCDM dilemma is that best alternatives are looked for among a collection which usually includes non-comparable pairs. Such non-comparable alternatives are often made comparable via a scalarization: the weighted sum method is such a (linear) scalarization but problems occur if the weights are not known (e.g., the ’unknown weight scenario’ in [14, p. 72]) or this method is not desirable at all. Here, a new ranking (a.k.a., scalarization) method is proposed which takes into account a whole bunch of predefined linear scalarizations at once and is monotone with respect to a vector preoder generated by a convex cone. This method can also detect certain elements of the Pareto frontier which do not necessarily lie on the boundary of the convex hull of the set of alternatives.
@@
A higher value of the ranking functions is ”better” if the goal is to maximize with respect to the underlying vector order $\leq_{C}$; a lower value should be considered better if the goal is minimization. In this paper, ”better” is associated with a higher value of the rank function, but this of course does not restrict generality. We also say that an alternative $z\in X$ dominates another $y\in X$ if $y\leq_{C}z$.
An easy way to find such an $r$ which is still widely used is the weighted sum scalarization: take $w\in C^{+}=\{v\in\mathrm{I\negthinspace R}^{d}\mid\forall z\in C\colon v^{\top}z\geq 0\}$ ($C^{+}$ is called the dual of the cone $C$) and define $r(z)=w^{T}z$. This makes the ranking a very subjective procedure: changing the weight vector even slightly might result in a different ranking. Moreover, different decision makers (e.g., reviewers of a project) might have different weight vectors. Therefore, many other methods have been suggested such as TOPSIS, ELECTRE, AHP, PROMETHEE etc.
@@
It originates from multivariate statistics where cone distribution functions were used to define quantiles of multidimensional random variables [6]. It will be shown that these functions can be used to define rankings (this idea is due to [9]) which enjoy very special rank reversal features. These features turn out to be useful properties instead of a nuisance since the situations in which a rank reversal occurs can be tracked and explained.
@@ 2 Cone ranking functions
The following definition introduces the basic concept of the paper.
@@
Proof. Assume $y\leq_{C}z$. Then $w^{T}y\leq w^{\top}z$ for all $w\in C^{+}$. Hence $x\in y-H^{+}(w)$ implies $x\in z-H^{+}(w)$. This implies $r_{X,w}(y)\leq r_{X,w}(z)$ for all $w\in C^{+}$ and consequently $r_{X,C}(y)\leq r_{X,C}(z)$. If $y<_{C}z$, then $y\in z-{\rm int\,}H^{+}(w)$ for all $w\in C^{+}\backslash\{0\}$, hence $z\not\in y-H^{+}(w)$. On the other hand $x\in y-H^{+}(w)\subseteq z-H^{+}(w)$. Hence $r_{X,w}(y)+1\leq r_{X,w}(z)$ for all $w\in C^{+}\backslash\{0\}$ which gives the result. The straightforward proof of the affine equivariance property (2.3) can be found in [6]. The last claim follows from the definition of the ranking functions $r_{X,C}$, $r_{X,D}$ together with a well-known property of polar cones: if $C\subseteq D$, then $D^{+}\subseteq C^{+}$.
Interpretation. The number $r_{w}(z)$ is precisely the number of alternatives which have a lower or the same weighted sum with weight distribution $w$ than $z$; the minimum over these numbers over all possible weight vectors gives $r_{X,C}(z)$. This means that no matter which feasible weight vector is chosen, $z$ has a higher weighted sum than at least $r_{X,C}(z)$ alternatives. If the task is to find an alternative which is maximal with respect to $\leq_{C}$, then one would look for $x\in X$ with $r_{X,C}(x)$ as large as possible–$x$ dominates as many points as possible no matter which weighted sum is chosen. This makes it clear that a ranking via $r_{w}$ or $r_{C}$ is a relative one: there is no ”objective” scale, the alternatives are only mutually compared, not with respect to an outside scale (such as temperature, for instance). Such a ranking is desirable for example in cases of indices for countries, economies, projects, candidates etc. where choices can be made only among a pool of available alternatives and one wishes to select the relative best. Roughly speaking, the function $r_{X,C}$ ranks an alternative higher if the minimal number of alternatives it dominates with respect to a feasible weight distribution is higher.
@@
One might call an alternative an outlier if it is not comparable to (and thus not dominated by) many others and has a very low $r_{X,C}$-value compared to the best ranked alternative–it has very different (not necessarily worse) features than the rest. It could be useful to identify such alternatives, e.g., for recommender systems since it could make sense to mix in such an alternative sometimes as a recommendation to provide options outside the usual ”bubble.”
@@
Figure 2.1 shows a situation where points with a maximal value of $r_{X,C}$ are not necessarily on the ”convex part” of the Pareto frontier, i.e., on the Pareto frontier of the convex hull of the points representing the alternatives. First, consider only the three black points: they all have rank 1 and are Pareto maximal with respect to the order generated by $C=\mathrm{I\negthinspace R}^{2}_{+}$. If one adds the three yellow points, then the three black ones are still Pareto maximal and the one in the middle gets the maximal ranking which now is 4. Clearly, this point does not belong to the convex hull of the now 6 data points. By the way of conclusion, this example shows that the ranking function $r_{X,C}$ can detect maximal alternatives not belonging to the ”convex part” of the Pareto frontier–even though only linear scalarizations enter the definition in (2.2). On the other hand, it also shows that the rank of a point particularly depends on how many other points it dominats with respect to the order generated by the cone: if one placed the three yellow points in Figure 2.1 close to the upper left or the lower right black point, one could generate maximal ranking for either of them. This feature will be further discussed in the sequel.
@@ 3 Rank reversal for cone ranking functions
Understanding this rank reversal feature contributes to understanding the proposed ranking method. First, a high ranking $r_{X,C}(y)$ compared to $r_{X,C}(x)$ does not mean that alternative $y$ is necessarily much better than $x$. It only means that $y$ dominates a higher number of alternatives if the ”worst” weight distributions are chosen, i.e., the $w$’s which provides the minimum in (2.2) for $x$ and $y$, respectively. Still, $x$ and $y$ can be non-comparable with respect to the original order $\leq_{C}$. In such a case, this could indicate that $x$ has very different features compared to $y$ and all alternatives dominated by $y$ (the yellow points in Figure 3.1). One might say that in such a case $y$ is a ”common” and $x$ is a ”rare” alternative and one would need to decide between a safe (common) and an adventureous (rare) option which are not comparable. To test the occurence of this feature, one can rank the alternatives according to $r_{X,C}$, then remove all alternatives which are dominated by the best one(s) and repeat the ranking with the reduced set. In the example of Figure 3.1, one would remove the yellow points after ranking all yellow/black points and get the upper right black point as the ”rare” option. Note, however, that very different ranks $r_{X,C}(x)$, $r_{X,C}(x)$ can also occur in cases where $x,y\in X$ dominate the same number of other alternatives.
@@
4 An example
The cone ranking function can produce results which are quite different from those generated by standard MCDM tools such as TOPSIS. The ranking of a student cohort is discussed as an illustrating example. The two criteria ”average mark in exams” and ”credit points achieved in a given time interval” are used. In the pictures, higher ranked individuals appear in lighter colours. Figure 4.1 shows the ranking obtained with TOPSIS: this method prefers–a little counterintuitively–the credit point criterion over the other. Figure 4.2 shows the result according to the cone ranking function with $C=\mathrm{I\negthinspace R}^{2}_{+}$: it gives higher rankings to alternatives in the upper right area.
@@ 4 An example
A similar example has been discussed in [7, Example 6.3] from a statistical point of view.
@@ 5 Discussion of the cone
In the examples above, the ”MCDM cone” $C=\mathrm{I\negthinspace R}^{d}_{+}$ was used. However, different cones are sometimes of advantage. The elements of the dual cone $C^{+}$ in Definition 2.1 can be seen as potential weight vectors for the $d$ different criteria. One may also observe that the condition $x\in z-H^{+}(w)$ in (2.1), (2.2) means $w^{T}(x-z)\leq 0$, thus it is positively homogeneous in $w$. Therefore, one can restrict the set of $w$’s to $B^{+}=\{w\in\mathrm{I\negthinspace R}^{d}_{+}\mid w_{1}+\ldots+w_{d}=1\}$ if $C=\mathrm{I\negthinspace R}^{d}_{+}$. The set $B^{+}$ includes all potential weight vectors for the $d$ criteria. If the decision maker wants to make sure that each criterion is given a minimal and a maximal weight, say $0\leq w_{i}^{min}\leq 1$ and $0\leq w_{i}^{max}\leq 1$ for $i\in\{1,\ldots,d\}$, respectively, then one can consider the set
@@
Moreover, this procedures gives more flexibility to the decision maker since the weight distribution does not have to be fixed in advance: the cone ranking function is a worst case ranking with respect to a variety of weight distributions. It is also useful to consider the minimizer(s) in (2.2): these are the weight vectors which gives the worst $w$-ranking of an alternative. Thus, if the decision maker has a preferred weight distribution (very) different from the minimizers in $r_{X,C}(x)$ for $x\in X$, then corresponding $w$-ranking of $x$ might even be much better than its $C$-ranking.
@@
6 Supervised classification models

### 14 Jan 2024 v2

@@ Abstract
Cone distribution functions from statistics are turned into Multi-Criteria Decision Making tools. It is demonstrated that this procedure can be considered as an upgrade of the weighted sum scalarization insofar as it absorbs a whole collection of weighted sum scalarizations at once instead of fixing a particular one in advance. As examples show, this type of scalarization–in contrast to a pure weighted sum scalarization–is also able to detect “non-convex” parts of the Pareto frontier. Situations are characterized in which different types of rank reversal occur, and it is explained why this might even be useful for analyzing the ranking procedure. The ranking functions are then extended to sets providing unary indicators for set preferences which establishes, for the first time, the link between set optimization methods and set-based multi-objective optimization. A potential application in machine learning is outlined.
Keywords. multi-objective optimization, ranking function, cone distribution function, rank reversal, set optimization, supervised learning
Mathematical Subject Classification. 68T05, 68Q32, 90C29
@@ 1 The ranking problem
The Multi-Criteria Decision Making (MCDM) dilemma is that best alternatives are looked for among a collection which usually includes non-comparable pairs. Such non-comparable alternatives are often made comparable via a scalarization: the weighted sum method is one of the simplest but problems occur if the weights are not known (e.g., the ‘unknown weight scenario’ in [20, p. 72]) or this method is not desirable at all. A new ranking (a.k.a., scalarization) method is proposed in this paper which takes into account a whole bunch of predefined linear scalarizations at once and is monotone with respect to a vector preoder generated by a convex cone. This method can also detect certain elements of the Pareto frontier which do not necessarily lie on the boundary of the convex hull of the set of alternatives.
@@
A higher value of the ranking functions is “better” if the goal is to maximize with respect to the underlying vector order $\leq_{C}$; a lower value should be considered better if the goal is minimization. In this paper, “better” is associated with a greater value of the rank function and therefore we also say that an alternative $z\in X$ dominates another $y\in X$ if $y\leq_{C}z$. Of course, this does not restrict generality.
The weighted sum scalarization is an easy way to find an $r$ satisfying (1.1): take $w\in C^{+}=\{v\in\mathrm{I\negthinspace R}^{d}\mid\forall z\in C\colon v^{\top}z\geq 0\}$ ($C^{+}$ is called the dual of the cone $C$) and define $r(z)=w^{T}z$. This makes the ranking a very subjective procedure: changing the weight vector even slightly might result in a different ranking. Moreover, different decision makers (e.g., reviewers of a project) might have different weight vectors. Therefore, many other methods have been suggested such as TOPSIS, ELECTRE, AHP, PROMETHEE etc.
@@
It originates from multivariate statistics where cone distribution functions were used to define quantiles of multi-dimensional random variables [10]. It will be shown that these functions can be used to define rankings (this idea is due to [15]) which enjoy very special rank reversal features. These features turn out to be useful properties instead of a nuisance since the situations in which a rank reversal occurs can be tracked and explained.
@@ 2 Cone ranking functions
The following definition introduces the basic concept of the paper. If $A$ is a finite set, #$\#A$ denotes its cardinality, i.e., the number of elements in $A$.
@@
Proof. Assume $y\leq_{C}z$. Then $w^{T}y\leq w^{\top}z$ for all $w\in C^{+}$. Hence $x\in y-H^{+}(w)$ implies $x\in z-H^{+}(w)$. This implies $r_{X,w}(y)\leq r_{X,w}(z)$ for all $w\in C^{+}$ and consequently $r_{X,C}(y)\leq r_{X,C}(z)$. If $y<_{C}z$, then $y\in z-{\rm int\,}H^{+}(w)$ for all $w\in C^{+}\backslash\{0\}$, hence $z\not\in y-H^{+}(w)$. On the other hand $x\in y-H^{+}(w)\subseteq z-H^{+}(w)$. Hence $r_{X,w}(y)+1\leq r_{X,w}(z)$ for all $w\in C^{+}\backslash\{0\}$ which gives the result. The straightforward proof of the affine equivariance property (2.3) can be found in [10]. The last claim follows from the definition of the ranking functions $r_{X,C}$, $r_{X,D}$ together with a well-known property of dual cones: if $C\subseteq D$, then $D^{+}\subseteq C^{+}$.
Interpretation. The number $r_{w}(z)$ is the number of alternatives which have a lower or the same weighted sum with weight distribution $w$ than $z$; the minimum over these numbers over all reasonable weight vectors gives $r_{X,C}(z)$. This means that no matter which feasible weight vector is chosen, $z$ has a greater weighted sum than at least $r_{X,C}(z)$ alternatives. If the task is to find an alternative which is maximal with respect to $\leq_{C}$, then one would look for $x\in X$ with $r_{X,C}(x)$ as large as possible–$x$ dominates as many points as possible no matter which weighted sum is chosen. This makes it clear that a ranking via $r_{w}$ or $r_{C}$ is a relative one: there is no objective scale, the alternatives are only mutually compared, not with respect to an outside scale (such as temperature, for instance). Roughly speaking, the function $r_{X,C}$ ranks an alternative higher if the minimal number of alternatives it dominates with respect to a feasible weight distribution is greater. Such a ranking is desirable for example in cases of indices for countries, economies, projects, candidates etc. where choices can be made only among a pool of available alternatives and one wishes to select the relative best.
@@
One might call an alternative an outlier if it is not comparable to (and thus not dominated by) many others and has a very low $r_{X,C}$-value compared to the best ranked alternative–it has very different (not necessarily worse) features than the rest. It could be useful to identify such alternatives, e.g., for recommender systems since it could make sense to mix in such an alternative sometimes as a recommendation to provide options outside the usual “bubble.”
@@
Figure 2.1 shows a situation where points with a maximal value of $r_{X,C}$ are not necessarily on the “convex part” of the Pareto frontier, i.e., on the Pareto frontier of the convex hull of the points representing the alternatives. First, consider only the three black points: they all have rank 1 and are Pareto maximal with respect to the order generated by $C=\mathrm{I\negthinspace R}^{2}_{+}$. If one adds the three yellow points, then the three black ones are still Pareto maximal and the one in the middle gets the maximal ranking which now is 4. Clearly, this point does not belong to the convex hull of the now 6 data points. By the way of conclusion, this example shows that the ranking function $r_{X,C}$ can detect maximal alternatives not belonging to the “convex part” of the Pareto frontier–even though only linear scalarizations enter the definition in (2.2). On the other hand, it also shows that the rank of a point particularly depends on how many other points it dominats with respect to the order generated by the cone: if one placed the three yellow points in Figure 2.1 close to the upper left or the lower right black point, one could generate maximal ranking for either of them. This feature will be further discussed in the sequel.
@@ 3 Rank reversal for cone ranking functions
Understanding this rank reversal feature contributes to understanding the proposed ranking method. First, a high ranking $r_{X,C}(y)$ compared to $r_{X,C}(x)$ does not mean that alternative $y$ is necessarily much better than $x$. It only means that $y$ dominates a greater number of alternatives if the worst weight distributions are chosen, i.e., the $w$’s which provides the minimum in (2.2) for $x$ and $y$, respectively. Still, $x$ and $y$ can be non-comparable with respect to the original order $\leq_{C}$. In such a case, this could indicate that $x$ has very different features compared to $y$ and all alternatives dominated by $y$ (the yellow points in Figure 3.1). One might say that in such a case $y$ is a “common” and $x$ is a “rare” alternative and one would need to decide between a safe (common) and an adventureous (rare) option which are not comparable. To test the occurence of this feature, one can rank the alternatives according to $r_{X,C}$, then remove all alternatives which are dominated by the best one(s) and repeat the ranking with the reduced set. In the example of Figure 3.1, one would remove the yellow points after ranking all yellow/black points and get the upper right black point as the “rare” option. Note, however, that very different ranks $r_{X,C}(x)$, $r_{X,C}(x)$ can also occur in cases where $x,y\in X$ dominate the same number of other alternatives.
@@
The cone ranking function can produce results which are quite different from those generated by standard MCDM tools such as TOPSIS. The ranking of a student cohort is discussed as an illustrating example. The two criteria “average mark in exams” and “credit points achieved in a given time interval” are used. In the pictures, higher ranked individuals appear in lighter colours. Figure 3.3 shows the ranking obtained with TOPSIS: this method puts more weight–a little counterintuitively–to the credit point criterion. Figure 3.4 shows the result according to the cone ranking function with $C=\mathrm{I\negthinspace R}^{2}_{+}$: it gives higher rankings to alternatives in the upper right area.
@@
A similar example has been discussed in [11, Example 6.3] from a statistical point of view.
@@ 4 Discussion of the cone
In the examples above, the “MCDM cone” $C=\mathrm{I\negthinspace R}^{d}_{+}$ was used. However, different cones are sometimes of advantage. The elements of the dual cone $C^{+}$ in Definition 2.1 can be seen as potential weight vectors for the $d$ different criteria. One may also observe that the condition $x\in z-H^{+}(w)$ in (2.1), (2.2) means $w^{T}(x-z)\leq 0$, thus it is positively homogeneous in $w$. Therefore, one can restrict the set of $w$’s to $B^{+}=\{w\in\mathrm{I\negthinspace R}^{d}_{+}\mid w_{1}+\ldots+w_{d}=1\}$ if $C=\mathrm{I\negthinspace R}^{d}_{+}$. The set $B^{+}$ includes all potential weight vectors for the $d$ criteria. If the decision maker wants to make sure that each criterion is given a minimal and a maximal weight, say $0\leq w_{i}^{min}\leq 1$ and $0\leq w_{i}^{max}\leq 1$ for $i\in\{1,\ldots,d\}$, respectively, then one can consider the set
@@
Moreover, this procedures gives more flexibility to the decision maker since the weight distribution does not have to be fixed in advance: the cone ranking function is a worst case ranking with respect to a variety of weight distributions. It is also useful to consider the minimizer(s) in (2.2): these are the weight vectors which gives the worst $w$-ranking of an alternative. Thus, if the decision maker has a preferred weight distribution (very) different from the minimizers in $r_{X,C}(x)$ for $x\in X$, then the corresponding $w$-ranking of $x$ might even be much better than its $C$-ranking.
@@
5 Extension to rankings of sets
The comparison and ranking of sets as a problem was introduced to MCDM by Zitzler and Thiele in the highly cited paper [25]. It was motivated by the need to compare the output sets of different MCDM algorithms which include potentially non-dominated alternatives. In the follow-ups [24, 26], several order relations for sets can be found, especially the ”dominates” relation (along with strict and weak versions) which coincides with the following order relation for sets: a set $A\subseteq\mathrm{I\negthinspace R}^{d}$ dominates another $B\subseteq\mathrm{I\negthinspace R}^{d}$, written $A\succeq_{C}B$, if
$\forall b\in B,\;\exists a\in A\colon a\geq_{C}b.$
In the vector preorder setting, this is equivalent to $A-C\supseteq B$. This relation is one of two which form the building block for the theory of set optimization: compare [9] for surveys of the relevant literature. It has been known before, for example, in Economics: compare [13] for some historic comments. To the best of our knowledge, the Thiele-Zitzler approach was developed independently of the set optimization literature (and vice versa) which can also be concluded from [26]: in this reference, the twin relation to $\succeq_{C}$ can be found (compare [26, p. 61, Sect. C]) which is more suited for minimization.
The performance measure $\mathcal{C}$ in [25, Sect. III, (7)] gives the value $\mathcal{C}(A,B)=1$ precisely if $A\succeq_{C}B$ for $A,B\subseteq X$ since it is the number of alternatives in $B$ which are dominated by at least one alternative in $A$ relative to the number of alternatives in $B$.
On the other hand, set extensions of the lower cone distribution function from [10] were shown to be in the same relation with set-valued quantiles as the cumulative distribution function with quantile function for univariate random variables–they form a Galois connection (and are even inverse to each other if the underlying random variable is continuous). These set extensions are monotone with respect to the underlying set order relation. A parallel extension of the ranking functions from Definition 2.1 can be used to rank sets of alternatives instead of single alternatives.
In the following, the symbols $\mathcal{P}(X)$, $\mathcal{P}(\mathrm{I\negthinspace R}^{d})$ denote the power sets of $X$ and $\mathrm{I\negthinspace R}^{d}$, respectively–including the empty set in each case.
Definition 5.1
The functions $R^{\rm\triangledown}_{w},R^{\rm\triangledown}_{C}\colon\mathcal{P}(\mathrm{I\negthinspace R}^{d})\to\mathrm{I\negthinspace N}$ defined by
[…]
are called set $w$-ranking for $w\in C^{+}$ and set cone ranking function, respectively, generated by the set $X$.
Clearly, if $A$ is a finite set–especially if $A\in\mathcal{\mathcal{missing}}P(X)$–then the suprema in Definition 5.1 can be replaced by maxima. Moreover, one of course has $R^{\rm\triangledown}_{w}(\{a\})=r_{X,w}(a)$ for singleton sets $A=\{a\}$.
Proposition 5.2
Both $R^{\rm\triangledown}_{w}$ and $R^{\rm\triangledown}_{C}$ are monotone with respect to $\succeq_{C}$.
Proof. First, we show $R^{\rm\triangledown}_{C}(A-C)=R^{\rm\triangledown}_{C}(A)$. Indeed, one has $a-z\leq_{C}a$ for all $a\in A$, $z\in C$, hence
$R^{\rm\triangledown}_{C}(A-C)=\sup_{a\in A,z\in C}r_{X,C}(a-z)\leq\sup_{a\in A}r_{X,C}(a)=R^{\rm\triangledown}_{C}(A)$
by monotonicity of $r_{X,C}$. On the other hand, $A-C\supseteq A$ implies
$R^{\rm\triangledown}_{C}(A-C)=\sup_{y\in A-C}r_{X,C}(y)\geq\sup_{a\in A}r_{X,C}(a)=R^{\rm\triangledown}_{C}(A),$
hence equality holds true. Next, take $A,B\subseteq\mathrm{I\negthinspace R}^{d}$ with $B\subseteq A-C$. Then
$R^{\rm\triangledown}_{C}(B)\leq R^{\rm\triangledown}_{C}(A-C)=R^{\rm\triangledown}_{C}(A)$
which proves the claim. The special case $C=H^{+}(w)$ yields the same result for $R^{\rm\vartriangle}_{w}$.
Interpretation. The definition of $R^{\rm\triangledown}_{w}$, namely $R^{\rm\triangledown}_{w}(A)=\sup_{a\in A}\#\{x\in X\mid w^{\top}x\leq w^{\top}a\}$, shows that $R^{\rm\triangledown}_{w}(A)$ asks to look for $a\in A$ which gives the maximal number of points in $X$ which have a weighted sum less than the one of $a$ with weight vector $w$. Consequently, $R^{\rm\triangledown}_{C}(A)=\sup_{a\in A}\min_{w\in C^{+}}\#\{x\in X\mid x\in a-H^{+}(w)\}$ ask for elements $a\in A$ for which the minimal number of points in $X$ with weighted sum less than the one for $a$ is maximal. Thus, it can be understood as a max-min performance measure for sets $A$: selecting a set is done by looking for best alternatives in sets with respect to $r_{X,C}$.
The comparison with the performance measure $\mathcal{C}$ from [25, Sect. III, (7)] cannot be made directly since the latter is a function of two sets. However, if one chooses the set $X$ of all $N$ alternatives as one, then the function on $\mathcal{P}(\mathrm{I\negthinspace R}^{d})$ defined by
$\mathcal{C}_{X}(A)=N\cdot\mathcal{C}(A,X)=\#\left\{x\in X\mid\exists a\in A\colon a\geq_{C}x\right\}$
is monotone with respect to $\succeq_{C}$ as one may easily check. This raises the question what the relationship between $R^{\rm\triangledown}_{C}$ and $\mathcal{C}_{X}$ is.
If $A$ includes all non-dominated elements of $X$, then $\mathcal{C}_{X}(A)=N$ since $X$ has only finitely many elements and thus enjoys the upper domination property, i.e., each element of $X$ is dominated by a non-dominated element. Moreover, $\mathcal{C}_{X}(X)=N$. Therefore, $\mathcal{C}_{X}(A)/N=\mathcal{C}(A,X)$ for $A\subseteq X$ can be understood as a measure for the performance of $A$ relative to $X$.
Potential outcomes. First, one always has $R^{\rm\triangledown}_{C}(\{z\})\geq\mathcal{C}_{X}(\{z\})$ for all singleton sets with $z\in\mathrm{I\negthinspace R}^{d}$. This is no longer true for non-singleton sets (compare the following examples).
Moreover, the ranking of sets via $R^{\rm\triangledown}_{C}$ can be different from the one via $\mathcal{C}_{X}$. Figure 5.1 shows three examples where in all cases the sets $A^{i}=\{x^{1},x^{2},x^{3}\}\backslash\{x^{i}\}$, $i=1,2,3$, are compared. Depending which points are added to the set $X$ the rankings of $A^{1},A^{2},A^{3}$ change.
First, the left picture in Figure 5.1 with no additional points gives $\mathcal{C}_{X}(A^{i})=2$ for $i=1,2,3$, i.e., $\mathcal{C}_{X}$ ranks all three sets as equal. On the other hand, $r_{X,C}(x^{1})=1$, $r_{X,C}(x^{2})=2$, $r_{X,C}(x^{3})=1$ and thus $R^{\rm\triangledown}_{C}(A^{1})=\max\{2,1\}=2$, $R^{\rm\triangledown}_{C}(A^{2})=\max\{1,1\}=1$, $R^{\rm\triangledown}_{C}(A^{3})=\max\{1,2\}=2$ which means that $A^{2}$ is ranked lower than the sets $A^{1}$, $A^{3}$ which are ranked equal. Since in all cases the cone is $C=\mathrm{I\negthinspace R}^{2}_{+}$ one might say that $R^{\rm\triangledown}_{C}$ gives a ranking which ranks sets higher if they include “compromise” non-dominated points such as $x^{2}$ in this example.
Secondly, the center picture of Figure 5.1 shows a situation where three dominated points are added which are not comparable with each other. This changes the $\mathcal{C}_{X}$-ranking to $\mathcal{C}_{X}(A^{1})=3$, $\mathcal{C}_{X}(A^{2})=4$ and $\mathcal{C}_{X}(A^{3})=5$. On the other hand, one has $r_{X,C}(x^{1})=3$, $r_{X,C}(x^{2})=3$, $r_{X,C}(x^{3})=1$ and thus $R^{\rm\triangledown}_{C}(A^{1})=\max\{3,1\}=3$, $R^{\rm\triangledown}_{C}(A^{2})=\max\{3,1\}=3$, $R^{\rm\triangledown}_{C}(A^{3})=\max\{3,3\}=3$. The $R^{\rm\triangledown}_{C}$-ranking is the same for all three sets while the $\mathcal{C}_{X}$-ranking is different.
Figure 5.1: Set ranking
Thirdly, in the right picture of Figure 5.1 two more points are added which are not comparable to any of the points in the previous situation. This does not change the $\mathcal{C}_{X}$-ranking compared to the center picture. On the other hand, $r_{X,C}(x^{1})=3$, $r_{X,C}(x^{2})=5$, $r_{X,C}(x^{3})=1$ and thus $R^{\rm\triangledown}_{C}(A^{1})=\max\{5,1\}=5$, $R^{\rm\triangledown}_{C}(A^{2})=\max\{3,1\}=3$, $R^{\rm\triangledown}_{C}(A^{3})=\max\{3,5\}=5$ which gives a different $R^{\rm\triangledown}_{C}$-ranking compared to the previous case.
In particular, these example also show that a set can be ranked higher with respect to the $R^{\rm\triangledown}_{C}$-ranking and lower with respect to the $\mathcal{C}_{X}$-ranking as well as vice versa.
The comparison of Figures 5.1 and 5.2 shows that rank reversal can occur with respect to the $R^{\rm\triangledown}_{C}$-ranking even if only alternatives are added which are non-comparable to all existing ones. This is inline with the correspondding feature of the point ranking via $r_{X,C}$: adding the red points in Figure 5.2 to the situation in the right picture of Figure 5.1 produces $R^{\rm\triangledown}_{C}(A^{1})=\max\{5,6\}=6$, $R^{\rm\triangledown}_{C}(A^{2})=\max\{3,6\}=6$, $R^{\rm\triangledown}_{C}(A^{3})=\max\{3,5\}=5$ which reverses the ranking of $A^{2}$ and $A^{3}$ in these two situations.
The addition of alternatives which are comparable to some existing ones can also lead to rank reversal with respect to $\mathcal{C}_{X}$ whereas no rank reversal can occur when only alternatives are added which are not comparable to the existing ones.
Figure 5.2: Set rank reversal
This section is concluded with a discussion of $\mathcal{C}_{X}$, $R^{\rm\triangledown}_{C}$ as unary indicators in the terminology of [26]. It follows directly from [26, Thm. 3.1] and monotonicity that both $\mathcal{C}_{X}$ and $R^{\rm\triangledown}_{C}$ generate weak refinements of the order relation $\succeq_{C}$ by
$A\succeq_{\mathcal{C}_{X}}B\quad$ $:\Leftrightarrow\quad\mathcal{C}_{X}(A)\geq\mathcal{C}_{X}(B)$
$A\succeq_{R^{\rm\triangledown}_{C}}B\quad$ $:\Leftrightarrow\quad R^{\rm\triangledown}_{C}(A)\geq R^{\rm\triangledown}_{C}(B).$
It remains to be clarified if these indicators are also refinements, i.e., if $A\succeq_{C}B$, $B\not\succeq_{C}A$ implies $\mathcal{C}_{X}(A)>\mathcal{C}_{X}(B)$ and parallel for $R^{\rm\triangledown}_{C}$. For definitions, compare [26, Def. 2.4, 2.5 and Thm. 3.1].
Proposition 5.3
The relation $\succeq_{\mathcal{C}_{X}}$ is a refinement for $\succeq_{C}$ on $\mathcal{P}(X)$. The relation $\succeq_{R^{\rm\triangledown}_{C}}$ is a weak refinement for $\succeq_{C}$ on $\mathcal{P}(X)$, but not a refinement in general.
Proof. If $A,B\subseteq X$, then $A\succeq_{C}B$ implies $\{x\in X\mid\exists a\in A\colon a\geq_{C}x\}\supseteq\{x\in X\mid\exists b\in B\colon b\geq_{C}x\}$. On the other hand, $B\not\succeq_{C}A$ means that there is $\bar{a}\in A$ such that $b\not\geq_{C}\bar{a}$ for all $b\in B$. Thus, one has $\bar{a}\in\{x\in X\mid\exists a\in A\colon a\geq_{C}x\}$ since $A\subseteq X$ and $\bar{a}\not\in\{x\in X\mid\exists b\in B\colon b\geq_{C}x\}$ which means $\mathcal{C}_{X}(A)\geq\mathcal{C}_{X}(B)+1$.
Already the situation depicted in the left graph of Figure 5.1 with $B=\{x^{2},x^{3}\}$ and $A=\{x^{1}\}\cup B$ shows that $A\succeq_{C}B$, $B\not\succeq_{C}A$ do not imply $R^{\rm\triangledown}_{C}(A)>R^{\rm\triangledown}_{C}(B)$.
The monotonicity property in Proposition 5.2 is stronger than the weak refinement property in [26, Def. 2.4, Thm. 3.1]. This shows that the relation $\succeq_{R^{\rm\triangledown}_{C}}$ is somehow in between a weak refinement and a refinement. The examples in Figures 5.1 and 5.2 also make it clear that $\succeq_{R^{\rm\triangledown}_{C}}$ reacts to the addition of non-comparable points while $\succeq_{\mathcal{C}_{X}}$ does not. Finally, let us mention that according to [26, Thm. 3.1] a Richter-Peleg utility representation of $\succeq_{C}$, a popular concepts in economics (see, for example, [2] and the references therein), is a refinement in the sense of [26, Def. 2.4]; apparently, also this connection has never been made before in the MCDM literature.
6 Classification models